Metallica Bearings, Inc., is a young start-up company. No dividends will be paid on the stock over the next 14 years because the firm needs to plow back its earnings to fuel growth. The company will pay a dividend of $6 per share 15 years from today and will increase the dividend by 7 percent per year thereafter.

  

If the required return on this stock is 13 percent, what is the current share price?

Multiple Choice

  • $17.16

  • $18.61

  • $18.97

  • $15.99

  • $18.07
Answer #1

Using DDM,

Share Price = D15/(R-G)

Price in year 14 = 6 /(0.13 - 0.07) = 100

Current Price = Price in year 14 / (1+required rate)^14

= 100 /(1+0.13)^14 = 18.07 Answer

The last option is correct.

Please let me know in case you have any queries and I will be happy to assist you.